# InstructLayout: Instruction-Driven 2D and 3D Layout Synthesis with Semantic Graph Prior

### v1

@@ 1 Introduction
[…]
[lines=2]Automatically synthesizing controllable and appealing 2D and 3D layouts has been a challenging task for computer vision and graphics [1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15], which requires building a generative model by multimodal conditional information. An ideal layout synthesis system should fulfill at least three objectives: (1) comprehending instructions in natural languages, thus providing an intuitive and user-friendly interface; (2) designing object compositions that exhibit aesthetic appeal and thematic harmony; (3) placing objects in appropriate positions adhering to their functions and regular arrangements.
To enhance the controllability of layout synthesis, we need to address natural language instructions as inputs. However, natural instructions for layout design often rely on abstract object relationships, which has posed a significant challenge for recent advancements in 2D and 3D layout synthesis. Previous works [16; 17; 18; 19; 20; 21] primarily exclude instructional prompts or do not prioritize them, leading to a low level of controllability and interpretability in the generated layouts. Other studies [22; 23; 24] utilize relation graphs to provide explicit control over object interactions, which are, however, too complicated and fussy for human users to specify. Moreover, previous works merely represent objects by categories  [22; 6; 19; 17] or low-dimensional features [25; 11; 12; 21] which lack visual appearance details, resulting in style inconsistency and constraining customization options in layout synthesis.
@@
To fit practical scenarios and promote the benchmarking of instruction-drive layout synthesis, we curate two high-quality datasets containing paired layouts and instructions with the help of large language and multimodal models [27; 28; 29] for both 2D E-commerce poster and 3D indoor scene synthesis tasks. Comprehensive quantitative evaluations reveal that InstructLayout surpasses previous state-of-the-art methods by a large margin in terms of both generation controllability and fidelity. Each essential component of our method is carefully verified through ablation studies.
@@ 2 Related Work
Graphs have been used to guide complex scene synthesis in the form of scene hierarchies [30; 31], parse trees [32], scene graphs [33; 34], etc. Wang et al. [25] utilize an image-based module and condition its outputs on the edges of a relation graph within each non-differentiable step. They also adopt an autoregressive model [35] to generate relation graphs, which are unconditional and have limited object attributes. Other works [22; 23; 24] adopt conditional VAEs [36; 37] with graph convolutional networks [38] to generate layouts. While offering high controllability, these methods demand the specification of elaborate graph conditions, which are notably more intricate than those driven by natural languages. With the advent of attention mechanisms [39], recent approaches [16; 17; 20; 21] can implicitly acquire object relations by self-attention and condition scene synthesis with texts by cross-attention. However, text prompts in these works tend to be relatively simple, containing only object categories or lacking layout descriptions, limiting the expressiveness and customization. Implicit relation modeling also significantly hinders their controllability.
@@
Numerous methods have been proposed to automatically synthesize 2D layouts for posters, user interfaces, E-commerce advertisements, etc. Prior research [7; 6; 40; 11; 19] has primarily concentrated on generating layouts from scratch on a blank canvas, while others [18; 41; 42; 10; 12] focus on image content-aware layout synthesis, leading to more practical applications. However, they are either confined to unconditional generation based solely on the canvas(blank or image prior) [7; 6; 41; 42] or relying on object categories [11; 19], lacking the capability to control the layout synthesis process using user-friendly instructions. Furthermore, due to the scarcity of high-quality datasets, only a few studies [12; 43] have delved into predicting 2D object attributes and their textual content, which is vital for generating realistic and informative layouts.
@@ 2.3 Generative Models for Graphs
There have been lots of endeavors on generative models for undirected graphs, molecules and scene graphs by autoregressive models [44; 45], VAEs [46; 47], GANs [48; 49] and diffusion models [50; 51; 52; 53]. However, none of them is capable of understanding text conditions. Longland et al. [54] and Lo et al. [55] employ VAE and GAN, respectively, for text-driven generation of simple undirected graphs without semantics. We present a pioneering effort to generate holistic semantic graphs with expressive instructions.
@@ 3.1 Problem Statement
Given instructions $\mathbf{y}$, our goal is to learn the conditional scene distribution $q(\mathcal{S}|\mathbf{y})$. Rather than direct modeling [17; 19; 12; 21], we employ well-structured and informative graphs to serve as general and semantic latent. Each graph $\mathcal{G}_{i}$ contains a node set $\mathcal{V}_{i}\coloneqq\{\mathbf{v}_{j}^{i}\}_{j=1}^{N_{i}}$ and a directed edge set $\mathcal{E}_{i}\coloneqq\{e_{jk}^{i}|\mathbf{v}_{j}^{i},\mathbf{v}_{k}^{i}\in\mathcal{V}_{i}\}$. A node $\mathbf{v}_{j}^{i}$ functions as a high-level representation of an object $\mathbf{o}_{j}^{i}$, and a directed edge $e_{jk}^{i}$ explicitly conveys the relations between objects.
@@ 3.2 Preliminary: General Diffusion Models
Diffusion models can accommodate both continuous and categorical distributions by a meticulous design of the diffusion kernel $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. A prominent choice for continuous variables is the Gaussian kernel [57; 58; 21]. Uniform transition is widely adopted for categorical variables [59; 60; 52]. In this work, we utilize both the discrete and continuous diffusion models for the two stages in InstructLayout respectively to model distinct layout attributes.
@@ 3.3.1 Feature Discretization
While for the 3D scene layout synthesis task, it’s hard to describe geometric shapes and visual appearance of 3D models by explicit categorical attributes, so we utilize a pretrained multimodal-aligned model OpenShape [26] to extract their semantic features. However, these high-dimensional features (e.g., $d=1280$ in OpenShape pointbert-vitg14) are too complicated to model. We circumvent this drawback by introducing a vector-quantized variational autoencoder (VQ-VAE) [62; 63] for feature vectors and quantize each of them by several learnable tokens from the maintained codebook, as depicted in Figure 2(a). The intuition behind it is that 3D models share general intrinsic characteristics, encompassing attributes like colors, materials, and basic geometric shapes. Indexing semantic features in a discrete space achieves a compact feature space for distribution modeling and dramatically reduces the cost of operating in a continuous space.
@@ 3.3.2 Discrete Semantic Graph Diffusion
After feature discretization, all attributes in a semantic graph are categorical: a graph $\mathcal{G}_{i}\coloneqq(\mathcal{C}_{i},\mathcal{F}_{i},\mathcal{E}_{i})$ with $N_{i}$ nodes, where $\mathcal{C}_{i}\coloneqq\{1,\dots,K_{c}\}^{N_{i}}$, $\mathcal{E}_{i}\coloneqq\{1,\dots,K_{e}\}^{N_{i}\times N_{i}}$ and $\mathcal{F}_{i}\coloneqq\{1,\dots,K_{f}\}^{N_{i}\times n_{f}}$, where $n_{f}$ is the number of discrete features. While it is possible to embed discrete variables in continuous spaces using one-hot encodings, it diminishes the sparsity inherent in the original data and imposes a substantial burden on network optimization. Instead, we propose to model the semantic graph prior by a mask-based discrete diffusion model.
For a scalar discrete random variable with $K$ categories $x\in\{1,\dots,K\}$, diffusion noise is defined by a series of transition matrices $\mathbf{Q}\in\mathbb{R}^{K\times K}$. The forward process at timestep $t$ is expressed as $q(x_{t}|x_{t-1})\coloneqq\mathbf{x}_{t}^{\top}\mathbf{Q}_{t}\mathbf{x}_{t-1}$, where $\mathbf{x}_{t}\in\mathbb{R}^{K}$ is the column one-hot encoding for $x_{t}$ and $[\mathbf{Q}_{t}]_{mn}\coloneqq q(x_{t}=m|x_{t-1}=n)$ is the probability that $x_{t-1}$ transits to the category $m$ from $n$. The probabilistic distribution of $x_{t}$ can be directly derived from $x_{0}$: $q(x_{t}|x_{0})\coloneqq\mathbf{x}_{t}^{\top}\bar{\mathbf{Q}}_{t}\mathbf{x}_{0}$, where $\bar{\mathbf{Q}}_{t}\coloneqq\mathbf{Q}_{t}\cdots\mathbf{Q}_{1}$.
Instead of commonly used Gaussian or uniform transitions for graph generation [50; 59; 51; 52], we propose to diffuse semantic graphs by independently masking graph attributes (i.e., object class $c$, discrete feature indices $f$ and relation $e$) by introducing an absorbing state [MASK] [60; 64] to each uniform transition matrix. For object class $c$, its transition matrix is defined as:
@@
by which $c_{t}$ has a probability of $\gamma_{t}^{c}$ to be masked, a probability of $\alpha_{t}^{c}$ to maintain the same, leaving a chance of $1-\gamma_{t}^{c}-\alpha_{t}^{c}$ for uniform sampling. [MASK] will always stay in its state. Transition matrices for $f$ and $e$, denoted as $\mathbf{Q}_{t}^{\mathbf{F}}$ and $\mathbf{Q}_{t}^{\mathbf{E}}$ respectively, exhibit analogous structures. Schedules of $(\alpha_{t},\beta_{t},\gamma_{t})$ are designed to admit that the initial states for semantic graphs are all masked.
Since the number of objects varies across different layouts, semantic graphs are padded by empty states to maintain a consistent number of $N$ objects. One-hot encodings for scalar variables $c$, $f$ and $e$ in a layout are denoted as $\mathbf{C}\in\mathbb{R}^{N\times(K_{c}+2)}$, $\mathbf{F}\in\mathbb{R}^{N\times n_{f}\times(K_{f}+2)}$ and $\mathbf{E}\in\mathbb{R}^{N\times N\times(K_{e}+2)}$ respectively. Here, “$+2$” accounts for each variable’s two extra states (i.e., empty state and mask state). A one-hot encoded semantic graph $G_{0}\coloneqq(\mathbf{C}_{0},\mathbf{F}_{0},\mathbf{E}_{0})$ at timestep $t$ is formulated as
@@
These benefits are critical, especially for intricate semantic graphs and diverse downstream generative tasks, compared with simple graph generative tasks [50; 51; 52]. An ablation study on the choice of $\mathbf{Q}$ is provided in Sec. 6.2.
@@ 3.4.1 Spatial Feature Generation
A continuous diffusion model with variance-preserving Gaussian kernels [57; 58] is adopted to learn $p_{\theta}(\mathbf{L}|\mathcal{G})$. Following Ho et al. [57], the variational bound in Equation 2 for the decoder $p_{\theta}(\mathbf{L}|\mathcal{G})$ is reweighted and simplified:
@@ 3.4.2 Tagline Generation for 2D Poster
After synthesizing semantic features $\mathbf{f}$ and spatial features $\mathbf{s}$, the pipeline for 3D scene synthesis is completed by retrieving 3D objects from an off-the-shelf dataset based on the generated $\mathbf{f}$ and transforming them to appropriate positions using the generated $\mathbf{s}$.
For 2D poster synthesis, we proceed to predict the textual contents in bounding boxes using all generated $\mathbf{f}$ and $\mathbf{s}$ associated with user-provided product descriptions and images. Concretely, we primarily adopt the method proposed by Gao et al. [43] to generate tagline captions. This method involves an autoregressive transformer decoder considering multimodal conditions such as background images, product descriptions, and spatial features. A progressive training strategy trains it through two tasks, including caption generation and caption matching. However, the absence of clean product images and semantic features of text bounding boxes results in suboptimal tagline generation. To address this issue, we utilize high-quality cleaned product images processed according to the strategy proposed in Sec. 5.1 and incorporate the generated semantic features from Sec 3.3 as an additional robust condition to guide generation.
@@ 3.5 Model Architecture
Graph Transformer. The semantic graph prior and spatial decoder share the same model architecture as shown in Figure 3(b). Since relation $e_{jk}$ can be determined by $e_{kj}$, only the upper triangular part of the relation matrix is necessary. Object categories and features together form input tokens for Transformers. Message passing on graphs is operated via node self-attention and node-edge fusion with FiLM [66], which linearly modulates edge embeddings and node attention matrices before softmax [67; 52]. Timestep for diffusion $t$ is injected by AdaLN [68; 69]. In the graph prior $p_{\phi}(\mathcal{G}|\mathbf{y})$, instructions and optional product images are embedded by a frozen multimodal encoder CLIP [70] and consistently influence network outputs by cross-attention mechanisms. Layout decoder $p_{\theta}(\mathcal{S}|\mathcal{G})$ is conditioned on semantic graphs by appending sampled Gaussian noises on node embeddings, which are then iteratively denoised to produce layout spatial features.
Permutation Non-invariance. Although $\mathcal{G}$ should ideally remain invariant to node permutations, invariant diffusion models could encounter learning challenges for multi-mode modeling. Thus, each node feature is added with positional encodings [39; 71; 21] before the permutation-equivariant Transformer. Exchangeability for graph prior distributions is strived by random permutation augmentation during the training process. Ablation on the permutation non-invariance is provided in Sec. 6.2.

### v3

@@ 1 Introduction
Automatically synthesizing controllable and appealing 2D and 3D layouts has been a challenging task for computer vision and graphics [1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14], which requires building a generative model by multimodal conditional information. An ideal layout synthesis system should fulfill at least three objectives: (1) comprehending instructions in natural languages, thus providing an intuitive and user-friendly interface; (2) designing object compositions that exhibit aesthetic appeal and thematic harmony; (3) placing objects in appropriate positions adhering to their functions and regular arrangements.
To enhance the controllability of layout synthesis, we need to address natural language instructions as inputs. However, natural instructions for layout design often rely on abstract object relationships, which has posed a significant challenge for recent advancements in 2D and 3D layout synthesis. Previous works [15; 16; 17; 18; 19; 20] primarily exclude instructional prompts or do not prioritize them, leading to a low level of controllability and interpretability in the generated layouts. Other studies [21; 22; 23] utilize relation graphs to provide explicit control over object interactions, which are, however, too complicated and fussy for human users to specify. Moreover, previous works merely represent objects by categories  [21; 4; 18; 16] or low-dimensional features [24; 9; 10; 20] which lack visual appearance details, resulting in style inconsistency and constraining customization options in layout synthesis.
@@
To fit practical scenarios and promote the benchmarking of instruction-drive layout synthesis, we curate two high-quality datasets containing paired layouts and instructions with the help of large language and multimodal models [26; 27; 28] for both 2D E-commerce poster and 3D indoor scene synthesis tasks. Comprehensive quantitative evaluations reveal that InstructLayout surpasses previous state-of-the-art methods by a large margin in terms of both generation controllability and fidelity. Each essential component of our method is carefully verified through ablation studies.
@@ 2 Related Work
Graphs have been used to guide complex scene synthesis in the form of scene hierarchies [29; 30], parse trees [31], scene graphs [32; 33], etc. Wang et al. [24] utilize an image-based module and condition its outputs on the edges of a relation graph within each non-differentiable step. They also adopt an autoregressive model [34] to generate relation graphs, which are unconditional and have limited object attributes. Other works [21; 22; 23] adopt conditional VAEs [35; 36] with graph convolutional networks [37] to generate layouts. While offering high controllability, these methods demand the specification of elaborate graph conditions, which are notably more intricate than those driven by natural languages. With the advent of attention mechanisms [38], recent approaches [15; 16; 19; 20] can implicitly acquire object relations by self-attention and condition scene synthesis with texts by cross-attention. However, text prompts in these works tend to be relatively simple, containing only object categories or lacking layout descriptions, limiting the expressiveness and customization. Implicit relation modeling also significantly hinders their controllability.
@@
Numerous methods have been proposed to automatically synthesize 2D layouts for posters, user interfaces, E-commerce advertisements, etc. Prior research [5; 4; 39; 9; 18] has primarily concentrated on generating layouts from scratch on a blank canvas, while others [17; 40; 41; 8; 10] focus on image content-aware layout synthesis, leading to more practical applications. However, they are either confined to unconditional generation based solely on the canvas(blank or image prior) [5; 4; 40; 41] or relying on object categories [9; 18], lacking the capability to control the layout synthesis process using user-friendly instructions. Furthermore, due to the scarcity of high-quality datasets, only a few studies [10; 42] have delved into predicting 2D object attributes and their textual content, which is vital for generating realistic and informative layouts.
@@ 2.3 Generative Models for Graphs
There have been lots of endeavors on generative models for undirected graphs, molecules and scene graphs by autoregressive models [43; 44], VAEs [45; 46], GANs [47; 48] and diffusion models [49; 50; 51; 52]. However, none of them is capable of understanding text conditions. Longland et al. [53] and Lo et al. [54] employ VAE and GAN, respectively, for text-driven generation of simple undirected graphs without semantics. We present a pioneering effort to generate holistic semantic graphs with expressive instructions.
@@ 3.1 Problem Statement
Given instructions $\mathbf{y}$, our goal is to learn the conditional scene distribution $q(\mathcal{S}|\mathbf{y})$. Rather than direct modeling [16; 18; 10; 20], we employ well-structured and informative graphs to serve as general and semantic latent. Each graph $\mathcal{G}_{i}$ contains a node set $\mathcal{V}_{i}\coloneqq\{\mathbf{v}_{j}^{i}\}_{j=1}^{N_{i}}$ and a directed edge set $\mathcal{E}_{i}\coloneqq\{e_{jk}^{i}|\mathbf{v}_{j}^{i},\mathbf{v}_{k}^{i}\in\mathcal{V}_{i}\}$. A node $\mathbf{v}_{j}^{i}$ functions as a high-level representation of an object $\mathbf{o}_{j}^{i}$, and a directed edge $e_{jk}^{i}$ explicitly conveys the relations between objects.
@@ 3.2 Preliminary: General Diffusion Models
Diffusion models can accommodate both continuous and categorical distributions by a meticulous design of the diffusion kernel $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. A prominent choice for continuous variables is the Gaussian kernel [56; 57; 20]. Uniform transition is widely adopted for categorical variables [58; 59; 51]. In this work, we utilize both the discrete and continuous diffusion models for the two stages in InstructLayout respectively to model distinct layout attributes.
@@ 3.3.1 Feature Discretization
While for the 3D scene layout synthesis task, it’s hard to describe geometric shapes and visual appearance of 3D models by explicit categorical attributes, so we utilize a pretrained multimodal-aligned model OpenShape [25] to extract their semantic features. However, these high-dimensional features (e.g., $d=1280$ in OpenShape pointbert-vitg14) are too complicated to model. We circumvent this drawback by introducing a vector-quantized variational autoencoder (VQ-VAE) [61; 62] for feature vectors and quantize each of them by several learnable tokens from the maintained codebook, as depicted in Figure 2(a). The intuition behind it is that 3D models share general intrinsic characteristics, encompassing attributes like colors, materials, and basic geometric shapes. Indexing semantic features in a discrete space achieves a compact feature space for distribution modeling and dramatically reduces the cost of operating in a continuous space.
@@ 3.3.2 Discrete Semantic Graph Diffusion
After feature discretization, all attributes in a semantic graph $\mathcal{G}_{i}\coloneqq(\mathcal{C}_{i},\mathcal{F}_{i},\mathcal{E}_{i})$ are categorical, where $\mathcal{C}_{i}\in\{1,\dots,K_{c}\}^{N_{i}}$, $\mathcal{F}_{i}\in\{1,\dots,K_{f}\}^{N_{i}\times n_{f}}$, and $\mathcal{E}_{i}\in\{1,\dots,K_{e}\}^{N_{i}\times N_{i}}$, where $n_{f}$ is the number of discrete features. While continuous embeddings such as one-hot vectors are commonly used, they often obscure the inherent sparsity of categorical data and complicate optimization. We instead model the semantic graph prior using a mask-based discrete diffusion process.
For a categorical variable $x\in\{1,\dots,K\}$, diffusion noise is defined by transition matrices $\mathbf{Q}t\in\mathbb{R}^{K\times K}$ over timesteps $t$. The forward process is given by $q(x_{t}|x_{t-1})\coloneqq\mathbf{x}_{t}^{\top}\mathbf{Q}_{t}\mathbf{x}_{t-1}$, where $\mathbf{x}_{t}$ is the column one-hot encoding of $x_{t}$ and $[\mathbf{Q}_{t}]_{mn}\coloneqq q(x_{t}=m|x_{t-1}=n)$ is the probability that $x_{t-1}$ transits to the category $m$ from $n$. The probabilistic distribution of $x_{t}$ can be directly derived from $x_{0}$: $q(x_{t}|x_{0})\coloneqq\mathbf{x}_{t}^{\top}\bar{\mathbf{Q}}_{t}\mathbf{x}_{0}$, where $\bar{\mathbf{Q}}_{t}\coloneqq\mathbf{Q}_{t}\cdots\mathbf{Q}_{1}$.
Instead of relying on commonly used Gaussian or uniform transitions for graph generation [49; 58; 50; 51], we adopt a diffusion process tailored for semantic graphs by independently masking graph attributes (i.e., object class $c$, discrete feature indices $f$, and relation $e$) by introducing an absorbing state, [MASK] [59; 63], into each uniform transition matrix. For the object class $c$, its transition matrix is defined as:
@@
by which $c_{t}$ has a probability of $\gamma_{t}^{c}$ to be masked, a probability of $\alpha_{t}^{c}$ to remain unchanged, and a remaining probability of $1-\gamma_{t}^{c}-\alpha_{t}^{c}$ to be uniformly resampled from the label space. The absorbing state [MASK] remains fixed once reached. Transition matrices for $f$ and $e$, denoted by $\mathbf{Q}_{t}^{\mathbf{F}}$ and $\mathbf{Q}_{t}^{\mathbf{E}}$ respectively, follow the same design. The schedules of $(\alpha_{t},\beta_{t},\gamma_{t})$ are crafted such that semantic graphs start fully masked at the beginning of the diffusion process.
Since the number of objects varies across different layouts, semantic graphs are padded by empty states to maintain a consistent number of $N$ objects. One-hot encodings for scalar variables $c$, $f$ and $e$ in a layout are denoted as $\mathbf{C}\in\mathbb{R}^{N\times(K_{c}+2)}$, $\mathbf{F}\in\mathbb{R}^{N\times n_{f}\times(K_{f}+2)}$ and $\mathbf{E}\in\mathbb{R}^{N\times N\times(K_{e}+2)}$ respectively. Here, “+2+2+ 2” accounts for each variable’s two extra states (i.e., empty state and mask state). A one-hot encoded semantic graph $G_{0}\coloneqq(\mathbf{C}_{0},\mathbf{F}_{0},\mathbf{E}_{0})$ at timestep $t$ is formulated as
@@
These benefits are critical, especially for intricate semantic graphs and diverse downstream generative tasks, compared with simple graph generative tasks [49; 50; 51]. An ablation study on the choice of $\mathbf{Q}$ is provided in Sec. 6.2.
@@ 3.4.1 Spatial Feature Generation
A continuous diffusion model with variance-preserving Gaussian kernels [56; 57] is adopted to learn $p_{\theta}(\mathbf{L}|\mathcal{G})$. Following Ho et al. [56], the variational bound in Equation 2 for the decoder $p_{\theta}(\mathbf{L}|\mathcal{G})$ is reweighted and simplified:
@@ 3.4.2 Tagline Generation for 2D Poster
After synthesizing semantic features $\mathbf{f}$ and spatial layouts $\mathbf{s}$, the pipeline for 3D scene synthesis is completed by retrieving 3D objects from an off-the-shelf dataset based on the generated $\mathbf{f}$ and transforming them to appropriate positions using the generated $\mathbf{s}$.
For 2D poster synthesis, we further generate textual taglines for each bounding box using all generated $\mathbf{f}$, $\mathbf{s}$ and user-provided product descriptions and images. We adopt the multimodal captioning framework proposed by Gao et al. [42], which uses an autoregressive Transformer decoder conditioned on background images, product descriptions and spatial layout. The model is trained progressively via caption generation and matching tasks.
However, the absence of clean product images and semantic features of text bounding boxes results in suboptimal tagline generation. To address this issue, we utilize high-quality cleaned product images processed according to the strategy proposed in Sec. 5.1 and incorporate the generated semantic features from Sec 3.3 as an additional conditioning. This results in more coherent and product-relevant tagline generation.
@@ 3.5 Model Architecture
Graph Transformer. The semantic graph prior and spatial decoder share the same model architecture as shown in Figure 3(b). Since relation $e_{jk}$ can be determined by $e_{kj}$, only the upper triangular part of the relation matrix is necessary. Object categories and features together form input tokens for Transformers. Message passing on graphs is operated via node self-attention and node-edge fusion with FiLM [65], which linearly modulates edge embeddings and node attention matrices before softmax [66; 51]. Timestep for diffusion $t$ is injected by AdaLN [67; 68]. In the graph prior $p_{\phi}(\mathcal{G}|\mathbf{y})$, instructions and optional product images are embedded by a frozen multimodal encoder CLIP [69] and consistently influence network outputs by cross-attention mechanisms. Layout decoder $p_{\theta}(\mathcal{S}|\mathcal{G})$ is conditioned on semantic graphs by appending sampled Gaussian noises on node embeddings, which are then iteratively denoised to produce layout spatial features.
Permutation Non-invariance. Although $\mathcal{G}$ should ideally remain invariant to node permutations, invariant diffusion models could encounter learning challenges for multi-mode modeling. Thus, each node feature is added with positional encodings [38; 70; 20] before the permutation-equivariant Transformer. Exchangeability for graph prior distributions is strived by random permutation augmentation during the training process. Ablation on the permutation non-invariance is provided in Sec. 6.2.